OpenStudy (dimples97):

Could someone check this answer for me please? If im right or wrong? 17. Find the roots of the quadratic equation x2 – 8x = 9 by completing the square. Show your work. x^2-8x=9 x^2-8x+(8/2)^2=9(8/2)^2 x^2-8x+16=25 √(x-4)^2= √25 x-4=5 x= 1, x= -9

OpenStudy (bunny12):

Everything looks right except your answer. I think you accidently messed up. 1-4 does not equal -5 and -9-4 does not equal 5. Do you know what you did wrong?

OpenStudy (bunny12):

x-4=5 or x-4=-5

OpenStudy (dimples97):

Oh geez.. i just confused myself sorry hahah

OpenStudy (dimples97):

I dont know what i just did can you help me answer it correctly please

OpenStudy (bunny12):

ok so when something is squared, you put a plus and a negative sign in front of it because 5^2=25 and (-5)^2=25

OpenStudy (bunny12):

so x-4=5 which is 5+4=9

OpenStudy (bunny12):

x-4=-5 is -5+4=-1

OpenStudy (bunny12):

Do you understand it was just a small mistake. :)

OpenStudy (bunny12):

Your answer was 1-4=5 but 1-4=-3 and -9-4=5 but it is actually -9-4=-13

OpenStudy (dimples97):

Ohhh okay I see, thankyou for checking (:

OpenStudy (bunny12):

No problem :)